The binomial coefficient of the 4th term in the expansion of (x - q)5 is
A. 15 B. 20 C. 10 D. 5

Grade:Upto college level

1 Answers

Latika Leekha
9 years ago
By binomial theorem, we know that
(ab)5 = a5 − 5a4b + 10a3b2 − 10a2b3 + 5ab4b5.
Hence, (x - q)5 = x5 − 5x4q + 10x3q2 − 10x2q3 + 5xq4q5.
Hence, the binomial coeficient of teh fourth term is -10.
Thanks & Regards
Latika Leekha
askIITians Faculty